OpenStudy (anonymous):

solve for x: sqrt(x+1) - 3x = 1

OpenStudy (anonymous):

Add 3x to both sides, then square both sides.

sam (.sam.):

sqrt(x+1) - 3x+3x = 1+3x sqrt(x+1) = 1+3x square both sides (x+1)=(1+3x)^2 expand and solve

OpenStudy (anonymous):

its 0. x=0